OpenStudy (anonymous):

Find a cubic function with the given zeros. 7, -3, 2

OpenStudy (anonymous):

answer choices f(x) = x3 - 6x2 - 13x - 42 f(x) = x3 - 6x2 + 13x + 42 f(x) = x3 - 6x2 - 13x + 42 f(x) = x3 + 6x2 - 13 + 42

OpenStudy (anonymous):

So \((x-(7))(x-(-3))(x-(2))\) = 0.